2.5 mL of a bacteria culture has been added to 5 mL of fresh media. What is the dilution factor for this media?
Answer #1

The dilution factor or the dilution is the initial volume divided by the final volume.

DF=Vi/Vf

Vi = 2.5 ml

Vf = 2.5 + 5 = 7.5 ml

DF=Vi/Vf = 2.5/7.5 = 0.33